For the problem below, 1 is an extraneous solution and the solution is 2 and 9. True or false? why?
square root of 2x+9-square root of x+1=2

Answer by edjones(8007)

sqrt(2x+9)-sqrt(x+1)=2
3x-(2sqrt(2x+9)*sqrt(x+1))+10=4 square each side
-(2sqrt(2x+9)*sqrt(x+1))=-3x-6
8x^2+44x+36=9x^2+36x+36 square each side
x^2-8x=0 Subtract the left side from the right.
x(x-8)=0
x=8, x=0
.
Ed